OpenStudy (anonymous):

give an example of an arithmetic sequence that is found in the real world. find the common difference and write a recursive and iterative rule for the sequence. same with a geometric sequence

OpenStudy (ybarrap):

\( Arithmetic ~Sequence \) Traveling on a road: 1km,2km, 3km, ... $$ a_1=1\\a_2=2\\a_3=3\\\cdots\\a_n=a_1+d(n-1) $$ Where d=1. Does this make sense?

OpenStudy (ybarrap):

\(Geometric~Sequence\) Monthly Bank Account (in $): $$\large{ month_1=1\\month_2=2\\month_3=4\\month_4=8\\ \cdots\\ month_n=2 \times month_{n-1} =2^{n-1} } $$ Any questions?

OpenStudy (anonymous):

thank you

OpenStudy (ybarrap):

yw